Please use the The Law of Total Proabilities to Solve. Please round to two decimal places.

Archie must travel to a foreign country where a de
Answer #1

From the law of total probability

\frac{2}{12}*0.4 + \frac{10}{12}*p=0.30

\frac{2*0.4 + 10*p}{12}=0.3

solving for p

p=\frac{0.3*12 -2*0.4}{10}=0.28

Therefore, probability that Archie will succumb to the flu and die while he is away = 0.28